OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. 1/4 + b = 10

OpenStudy (anonymous):

b = (10 - 1/4) b = (40/4 - 1/4) b = 39/4

OpenStudy (anonymous):

b = (10 - 1/4) b = (40/4 - 1/4) b = 39/4